OpenStudy (anonymous):

Solve. a + 7 = 1 a = A.8 B.–8 C.–6 D.6

OpenStudy (anonymous):

1 - 7 = ?

OpenStudy (anonymous):

6!

OpenStudy (anonymous):

Just subtract 7 from 1

OpenStudy (anonymous):

wrong @jinxhead20²

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

wow i feel dumb

OpenStudy (anonymous):

It's C !

OpenStudy (anonymous):

@tatyy

OpenStudy (anonymous):

OHHHHHHHHHHH

OpenStudy (anonymous):

C is correct.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

y + 16 = –2 A.18 B.14 C.–14 D.–18 do i do the same thing to this cause i just got all of them wrong but the one yall told me

OpenStudy (anonymous):

-2-16=?

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

what? im lost now

OpenStudy (anonymous):

answer is D.

OpenStudy (anonymous):

how did yall do that

OpenStudy (anonymous):

Subtracting from a negative really means you're adding to it.

OpenStudy (anonymous):

-2-16=? same thing @Novalynne said

OpenStudy (anonymous):

Ex. If you have six missing socks and you lose two more you now have 8 missing socks. -6-2=-8

OpenStudy (anonymous):

ohhhh thank you

OpenStudy (anonymous):

You're welcome!